Space Stations

NASA ISS On-Orbit Status 20 June 2014

By Marc Boucher
Status Report
June 24, 2014
Filed under , ,
NASA ISS On-Orbit Status 20 June 2014
European Space Agency astronaut Alexander Gerst, Expedition 40 flight engineer, is pictured floating freely in the Columbus laboratory of the International Space Station.

Today: Biological Rhythms: Flight Engineer (FE)-6 Gerst removed a digital walk holter and actiwatch to end a data take for Biological Rhythms and saved the data to a medical laptop. Biological Rhythms 48 hours studies the effects of long-term microgravity exposure on heart function by analyzing an astronaut’s electrocardiogram for 48 hours.
Capillary Flow Experiment (CFE) Operations: Commander (CDR) Swanson set up hardware on the Maintenance Work Area (MWA) and conducted an Interior Corner Flow (ICF) test using the ICF5 hardware which contains a square shaped rectangular section equally divided into four smaller sections. CFE ICF studies capillary flow in interior corners of vessels with various geometries. Structured inside tanks providing interior corners are used in the design of fuel tanks to promote fuel flow to the outlet of the tank in the absence of gravity. Experimental results will provide the necessary boundary conditions as a function of container cross section and fill fraction. The benchmarked theory can be used to improve propellant management aboard spacecraft.

Education Payloads Operations (EPO) Top Experiment: Gerst started recording a European Space Agency (ESA) educational video for Top Experiment. The video demonstrates the physics of soap bubbles in micro-gravity and will be used for educational purposes on the ground.

Burning and Suppression of Solids (BASS)-II Cable Troubleshoot: On Monday, the BASS monitors stopped displaying the experiment video and still camera images after completing four flame tests. This monitor is specific to the BASS investigation and displays fuel, air, and fan readings during test burns. On Thursday FE-5 Wiseman conducted troubleshooting of the BASS-II video cable connector by taking conductivity measurements of the video cable. Wiseman placed kapton tape around the y-branch of the cable but was not able to reproduce the previous conductivity measurements. Today Wiseman tested the cable by connecting one side to the Microgravity Science Glovebox (MSG) and connecting the three remaining connectors one by one to the still camera, experiment assembly and the video camera to see if the breaker tripped during any of these connections. The cable connections were successful, but the video was still intermittent. Wiseman created a sling using kapton tape to hold the cable in place. The kapton tape sling was successful and BASS-II operations are scheduled to resume next week.

Internal Wireless Integrated System (IWIS) Remote Sensor Unit (RSU) Firmware Reload: Wiseman reloaded RSUs in the Service Module (SM), Lab, Functional Cargo Block (FGB), Node 3 and Japanese Experiment Module (JEM). The reload was to recover from suspected firmware corruption that was preventing the RSUs from providing valid acceleration data.

Post Extravehicular Activity: Following yesterday’s Russian EVA, today Gerst transferred USOS tools and equipment from the Russian Segment (RS) to the USOS.

Today’s Planned Tasks

Laptop RSS1,2, Reboot / r/g 1009 [Complete]
Laptop RS1(2) Reboot / (Onboard Computer System [Complete]
SM ??? (Caution & Warning Panel) Test [Complete]
Daily Planning Conference (S-band) [Complete]
BLR48 – Memory Card Replacement. [Complete]
MSG – Glovebox Activation [Complete]
Removal of US TV cameras, Lights, and REBA power unit from Orlan-MK Suits [Complete]
On MCC Go ??? 427 and DC1 Interface Leak Check. [Complete]
CFE – Hardware Test Ops Review [Complete]
BLR48 – Data Save [Complete]
Post-EVA Orlan debrief with specialists [Complete]
Photo/TV Camcorder Setup Verification [Complete]
EPO Top Experiment – Unstow Hardware [Complete]
Pressurization of Elektron-VM Liquid Unit before Activation. [Complete]
Orlan Orlan No. 4, No. 5 Feedwater Bladder Refill [Complete]
US Tools and Equipment Handover to USOS from RS after EVA-38 [Complete]
DC1-?? Hatch Opening after EVA-38 [Complete]
Stow T2 Exercise Video Equipment [Complete]
Handover of Photo Cameras to USOS after EVA [Complete]
Food Frequency Questionnaire [Complete]
On MCC Go ??-???427 Hatch Opening after EVA-38 [Complete]
EVA Camera Reconfig [Deferred]
??? 427 (DC1) deactivation, air duct installation /IRS Activation/Deactivation [Complete]
CFE – Preparation and Hardware Setup [Complete]
Drying Orlan-MK Water Feed Line [Complete]
EPO Top Experiment – Equipment Setup [Complete]
Orlan-MK Drying – start [Complete]
EPO Top Experiment – Review Big Picture Words [Complete]
CFE-ICF5 – Experiment Ops [Complete]
EPO-TOP Experiment Ops [Complete]
On MCC Go ISS Repress by 10 mm using air from the 2nd section of ??? 427 (Aft) ???? (start) [Complete]
Post-EVA debrief. [Complete]
IWIS RSU firmware reload [Complete]
EPO-TOP Equipment Stowage [Deferred]
BASS troubleshooting [Complete]
??? Maintenance [Complete]
On MCC Go ISS Repress Using ??? 427 (Aft) ???? Section 2 (terminate) [Complete]
Pille Dosimeter Reading after RS EVA / r/g 5898 [Complete]
Orlan-MK Drying (switching to backup fan) [Complete]
On MCC Go De-installing ??? 427 (DC1) Docking Mechanism.r/g 4412 [Complete]
JRNL – Journal Entry [Complete]
IMS and Stowage Conference [Complete]
Capillary Flow Experiment Hardware Teardown [Complete]
Post-EVA Private medical conference [Complete]
HAM Radio Session [Complete]
Set up the first Orlan-MK battery pack 825?3 for discharge and start discharging the first pack
NANO – Module 39 Removal [Complete]
Winscat [Complete]
Disassembly and REMOTE RS LAPTOP stowage / r/g 5892 [Complete]
ECLSS Wastewater Storage Tank Assembly (WSTA) Fill [Complete]
Daily Planning Conference [Complete]
Orlan-MK Drying – terminate
Flight Director/ISS Crew Tagup [Complete]
Setting up Orlan and ??? for storage mode
MSG – Glovebox Powerdown [Complete]
Task List

ARED Cylinder Evac
Sabatier purge
IMS update
WHC KTO replace
EVA video download
Uragan photo
ECOK H/W reconfig

Nominal commanding
Three-Day Look Ahead:

Saturday, 06/21: Crew off duty
Sunday, 06/22: Crew off duty
Monday 06/23: ACE, ITCS Sampling, BASS
QUICK ISS Status – Environmental Control Group:

Component – Status
Elektron – On
Vozdukh – Manual
[???] 1 – SM Air Conditioner System (“SKV1”) – On
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Stop
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Standby
Urine Processing Assembly (UPA) – Norm
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full Up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.